OpenStudy (annalee607):

1)What is the area of a regular decagon (12 sides) with side length of 9 cm and apothem of 16.8 cm? Round to the nearest tenth. 2)What is the area of a regular Pentagon with a side of 10 in? Round to the nearest tenth.

Mehek (mehek14):

Pentagon: \(\dfrac{1}{4}\sqrt{5(5+2√5)}a^2\)

Mehek (mehek14):

and for #1, decagon is 10 sides A dodecagon is 12 sides so \(\dfrac{16.8*9}{2}*12\)

OpenStudy (annalee607):

What about this one. Two similar Trapezoids have areas of 432 cm(squared) and 48 cm(squared). What is the ratio of similarity?

Mehek (mehek14):

You can do 48:432 and then simplify